OpenStudy (anonymous):

What is the y-value for f(x) = tan(x) when x = -45°?

OpenStudy (dan815):

tan(x)=sin(x)/cos(x)

OpenStudy (dan815):

and sin(-45)=-sin(45) and cos(-45)=cos(45)

OpenStudy (dan815):

and sin(45)=cos(45), so tan(-45)=-1